# Pauli Stabilizer Models for Gapped Boundaries of Twisted Quantum Doubles and Applications to Composite Dimensional Codes

### v3

@@ Abstract
We provide new algorithms and provide example constructions of stabilizer models for the gapped boundaries, domain walls, and $0D$ defects of Abelian composite dimensional twisted quantum doubles. Using the physically intuitive concept of condensation, our algorithm explicitly describes how to construct the boundary and domain-wall stabilizers starting from the bulk model. This extends the utility of Pauli stabilizer models in describing non-translationally invariant topological orders with gapped boundaries. To highlight this utility, we provide a series of examples including a new family of quantum error-correcting codes where the double of $\mathbb{Z}_{4}$ is coupled to instances of the double semion (DS) phase. We discuss the codes’ utility in the burgeoning area of quantum error correction with an emphasis on the interplay between deconfined anyons, logical operators, error rates and decoding. We also augment our construction, built using algorithmic tools to describe the properties of explicit stabilizer layouts at the microscopic lattice-level, with dimensional counting arguments and macroscopic-level constructions building on pants decompositions. The latter outlines how such codes’ representation and design can be automated. Going beyond our worked out examples, we expect our explicit step-by-step algorithms to pave the path for new higher-algebraic-dimensional codes to be discovered and implemented in near-term architectures that take advantage of various hardware’s distinct strengths.
@@ I Introduction
As the number of qubits increases, it is becoming of growing importance to determine which quantum error-correcting codes are best suited to maximize performance on different quantum hardware architectures. In this work, we provide a new direction to help make this determination. We work within the context of the family of topological codes based on Abelian twisted quantum doubles [4, 17] which encompasses generalizations to the well-studied surface code. Here we explore the interplay of three generalizations to the surface code: local Hilbert space dimension, boundaries and defects, and spatial anisotropy to provide new routes towards error correction.
@@
As we will explore in this work, higher composite-dimensional local Hilbert spaces provide routes to utilize multiple topological phases simultaneously in one quantum error-correcting code. This direction is less explored since translationally invariant systems are simpler to study. Translationally invariant (generalized) Pauli stabilizer codes in prime dimensions were shown to be equivalent to copies of higher-dimension toric codes $D(\mathbb{Z}_{d})$. Additionally, for composite-dimension qudits, translationally invariant Pauli stabilizer models can realize the bulk of Abelian twisted quantum doubles [40]. These works of Refs. 41, 42, 43, 40, while covering all qudit dimensions, did not address the case of translationally non-invariant models. We here explore spatially anisotropic codes with an additional emphasis on the richness that composite-dimension qudits offer.
@@
However, constructing spatially anisotropic codes requires a careful treatment of the boundaries that separate distinct topological phases [44, 45, 46, 47, 48, 49, 50, 51, 52]. In particular, the existing abstract constructions [45, 47] of these boundaries and defects may not be experimentally efficient in certain cases. For example, while the bulk of each phase separately have a representation in terms of familiar Pauli stabilizer models, it is unclear if the gapped boundaries and defects can also have a Pauli stabilizer description. Answering this question has practical consequences for implementation as well as for assessing the computing power of these codes. A simple systematic method to construct these boundaries and 0D defects within the Pauli Stabilizers is still lacking.
@@
In this work, we show that boundaries, domain walls and point-like defects of Abelian twisted quantum doubles can be written in terms of generalized Pauli stabilizers, thus generalizing result of Ref. 40. By providing an algorithm to construct them starting from the bulk theory we prove this constructively. Our construction relies on local condensation, so it does not introduce new degrees of freedom [51, 52], non-Pauli stabilizers [45, 47], nor rely on lattice defects [53]. It also stays within the framework of topological phases without reverting to gauging symmetries of their underlying symmetry-protected phases [51]. It also automatically produces topologically complete models, meaning that all logical operators are macroscopic [54, 50]. Even other choices for the initial quantum double stabilizers, like spatially disconnected stabilizers [50], will automatically give models with topological order. As such, the resulting Hamiltonian for the boundary will also be described by local stabilizers. This construction offers both a simple intuitive understanding of the boundaries as well as an efficient method that can automate boundary construction for various phases and lattices.
@@
Our construction further facilitates the exploration of new family of codes that comprise of generalized holes, boundaries and defects of higher-dimensional codes. For simplicity, we exemplify this new family of codes by studying the smallest composite number $d=p\times q$ where $p=q=2$. This could be realized by four-dimensional qudits or, alternatively in the near term, by grouping qubits in pairs. We find new qualitative properties for higher dimensional codes such as tunable error biases which can be exploited to simplify noise models [55]. Their syndrome extraction process can also exhibit advantages compared to the qubit case.
@@ II Summary of main results
To begin, Sec. III highlights the power of our constructive algorithm by providing a variety of new quantum error correcting codes. Our results are based on the quantum double and twisted quantum double formalisms. Specifically, we work with the double of $\mathbb{Z}_{4}$, $D(\mathbb{Z}_{4})$, as a base model. Here we consider periodic and three types of distinct open boundaries. Further, we use our algorithm to derive a set of (internal) domain walls which generalize Bombin’s [53] $\mathbb{Z}_{2}$ twist. We describe how, with the combination of external boundaries and internal domain walls (co-dimension 1 defects) one realizes codes of logical dimension $\widetilde{D}=4^{M}$. This may be achieved, for example, as $M=0$ (planar or spherical), $M=1$ (surface code, not periodic), $M=2$ (surface or toric), or $M>2$ (appropriate $2M$-boundary surface or $M$-torus or other combinations involving twists) of logical $d=4$-qudits. This is then enriched by additionally including local patches (co-dimension 0 defects), of a Double Semion phase (DS) condensed from $\mathbb{Z}_{4}$. This generalizes the notion of punctures and adds, with $N+1$ DS patches, $N$ qubits for a total composite logical dimension of $\widetilde{D}=4^{M}\times 2^{N}$.
Building on these constructions and intuition, Sec. III.3 gives a high-level description of our family of new quantum error correcting codes. We provide one instance of the code that uses a parent $\mathbb{Z}_{4}$ phase with multiple localized regions (termed patches) of DS phases resulting in the $2^{N}$ contribution. We provide the logical Pauli operators that act on the patch codes’ qubit codespace. We then discuss how, in addition to combining logical qubits of different dimensions into a mutually compatible code, these qubits offer enhanced topological protection by virtue of the logical operators’ higher weight, which lowers error probabilities.
@@
After describing our new codes, Sec. IV presents the main systematic methodology which led to our codes’ discovery. Specifically, we provide an algorithm (1) that concretely illustrates how to gauge a parent code—meaning to insert boundaries, domain walls, and 0D defects into Abelian twisted quantum doubles—into the new codes presented. We emphasize the concept of anyon condensation, which is used as a tool to analyze high-level code properties and as a sub-algorithm. Condensation results in code gauging from the bottom up (i.e., microscopic $\rightarrow$ macroscopic) within a generalized Pauli stabilizer formalism which is gnostic to bulk and boundary degrees of freedom.
@@
To provide intuition and to see the algorithm in action, a series of boundaries for well-known codes are described within this formalism in Sec. IV.2. Domain walls are also constructed systematically in Sec. IV.4. The generalization from the $\mathbb{Z}_{2}$ twist to $\mathbb{Z}_{4}$ is exposed in Ex. IV.4. An important example of the domain wall between $\mathbb{Z}_{4}$ and DS is treated heuristically in Ex. IV.5 and, using the full machinery of our construction, in App. A. The construction of 0D defects is detailed in App. B.
@@
To study the properties of inserting these topological components into different topological codes under a unified framework, we begin in Sec. V which computes the ground state degeneracies in the presence of multiple boundaries, domain walls, and defects. These counting arguments are based on microscopic commuting projector stabilizer models. They are central to deriving the logical dimension of our family of new quantum error correcting codes. Ultimately, the properties of the topological codes do not depend on the exact layout and this enables us to discuss codes at scale.
@@
In Sec. VI, generic code description, that is agnostic to concrete stabilizer models, is used to compute the codespace dimension and provide intuition about the logical operators of different codes. This is carried out using the pants decomposition of any g-genus 2D orientable manifold in conjunction with the topological components of codes. We provide intuition by treating the $\mathbb{Z}_{4}$-DS patch code confirming results of Sec. V. Finally, Sec. VII concludes by discussing this work’s implications, open questions, and future research directions.
III The $\mathbb{Z}_{4}$ Family of Codes
@@
The bulk of each TQFT is described by a Hamiltonian $H_{A}$ for the Ath TQFT. In addition to the bulk of the topological phase, the phase may be equipped with external boundaries with the vacuum, as described by $H_{\partial A}$. Further, domain walls may separate two topological phases. The boundary between each pair ($A,B$) of TQFTs will be described by another Hamiltonian $H_{A\cap B}$. Gapped domain walls of two topologically non-trivial phases are closely related to their boundaries with the vacuum. In fact, as is discussed in Sec. IV.4, a gapped domain wall between two topological phases $A$ and $B$ is precisely the gapped boundary of the topological order $A\otimes B$ [45, 46]. If $B=A$, the domain wall may be seen as an internal domain wall of the phase A{A}. Altogether, the composite system will be described by a Hamiltonian
$H_{\text{tot}}$ $=\sum_{A}H_{A}+\sum_{B}H_{B}+\sum_{\partial A}H_{\partial A}+\sum_{\partial B}H_{\partial B}$ (1)
$\quad+\sum_{A,B}H_{A\cap B}$
In the lattice picture, the models are defined on a 2D orientable surface $\Sigma$, possibly containing a spatial boundary $\partial\Sigma$. The surface is then triangulated into a lattice $\Gamma$ consisting of vertices $V$, edges $E$, and plaquettes $P$ (faces). Each spatial boundary $\partial\Sigma$ has to be associated with a consistent boundary theory that is compatible with the bulk theory.
In (twisted) quantum double models, a qudit, taking values in a finite group $G$, is associated with each edge of the lattice $e\in E$ 111the terms edge and qudit are essentially synonymous in this work. These qudits live in a Hilbert space $\mathcal{H}_{e}$ spanned by the orthonormal basis generated by the group elements $\{|g\rangle_{e}:g\in G\}$. The total Hilbert space is then the tensor product of all the edge Hilbert spaces $\mathcal{H}_{tot}=\bigotimes_{e\in E}\mathcal{H}_{e}$. The case where the group is $\mathbb{Z}_{2}$ corresponds to the toric code, where a qubit is associated with each edge [4]. Since the basic results in this paper can also be shown in Abelian groups, we focused on the case of the cyclic group $\mathbb{Z}_{4}$, which has a simple implementation using two qubits or $d=4$ qudits (hereafter referred to as qudits). In addition, the Hamiltonian terms are built from the generalized Pauli stabilizers, defined below.
@@
With appropriate definitions, all physical properties of the model do not depend on the microscopic details of the lattice. Any triangulation of the surface $\Sigma$ will give rise to the same topological phase [4, 17]. For simplicity, we follow the most prominent conventions, as they appear in the literature, and work with a square lattice.
In Sec. III.1 we review the $\mathbb{Z}_{4}$ surface code, which is the generalization of the surface code for the case of qudits [4]. In addition, by deriving the gapped boundaries and domain walls, we use $\mathbb{Z}_{4}$ as a tool to review defects and edges. Afterwards, in Sec. III.2 we review a Doubled Semion code construction which is a twisted quantum double version of the $\mathbb{Z}_{2}$ toric code [17], and is also closely related to the $\mathbb{Z}_{4}$ code [40]. In Sec. III.3, we discuss hybrid codes of $\mathbb{Z}_{4}$ and DS and discuss their QECC properties.
III.1 Description of $\mathbb{Z}_{4}$

### v4

# Composite-dimensional Topological Codes with Boundaries and Defects

@@ Abstract
We introduce new algorithms and provide example constructions of stabilizer models for the gapped boundaries, domain walls, and $0D$ defects of Abelian composite-dimensional twisted quantum doubles. Using the physically intuitive concept of condensation, our algorithm explicitly describes how to construct the boundary and domain-wall stabilizers starting from the bulk model. This extends the utility of Pauli stabilizer models in describing non-translationally invariant topological orders with gapped boundaries. To highlight this utility, we provide a series of examples, including a new family of quantum error-correcting codes where the double of $\mathbb{Z}_{4}$ is coupled to instances of the double semion (DS) phase. We discuss the codes’ utility in the burgeoning area of quantum error correction with an emphasis on the interplay between deconfined anyons, logical operators, error rates, and decoding. We also augment our construction, built using algorithmic tools to describe the properties of explicit stabilizer layouts at the microscopic lattice-level, with dimensional counting arguments and macroscopic-level constructions building on pants decompositions. The latter outlines how such codes’ representation and design can be automated. Our results are validated by a series of error-correcting threshold calculations comparing our code’s performance with standard surface codes. To do so, we introduce a composite dimensional belief propagation decoder with ordered statistics that utilizes combination sweeps. Going beyond our worked-out examples, we expect our explicit step-by-step algorithms to pave the path for new higher-dimensional codes to be discovered and implemented in near-term architectures that take advantage of various hardware’s distinct strengths.
@@ I Introduction
As the number of qubits increases, it is of growing importance to determine which quantum error-correcting codes are best suited to maximize performance on different quantum hardware architectures. In this work, we provide a new direction to help make this determination. We work within the context of the family of topological codes based on Abelian twisted quantum doubles [107, 92], which encompasses generalizations to the well-studied surface code. Here, we explore the interplay of three generalizations to the surface code: local Hilbert space dimension, boundaries and defects, and spatial anisotropy to provide new routes towards error correction.
@@
As we will explore in this work, higher-dimensional composite local Hilbert spaces provide a route to simultaneously utilize multiple topological phases in a single quantum error-correcting code. This direction is less explored, since translationally invariant systems are simpler to study [84]. Translationally invariant (generalized) Pauli stabilizer codes in prime dimensions were shown to be equivalent to copies of higher-dimension toric codes $D(\mathbb{Z}_{d})$ [85]. Additionally, for composite-dimension qudits, translationally invariant Pauli stabilizer models can realize the bulk of Abelian twisted quantum doubles [66]. These works of Refs. 23, 25, 85, 66, while covering all qudit dimensions, did not address the case of translationally non-invariant models. We here explore spatially anisotropic codes with an additional emphasis on the richness that composite-dimension qudits offer. This spatial anisotropy is known to enable new properties like universality with non-Abelian islands [123].
@@
However, constructing spatially anisotropic codes requires a careful treatment of the boundaries that separate distinct topological phases [20, 12, 108, 52, 53, 112, 37, 130, 129, 128]. This was absent from previous works that considered Pauli stabilizer models of Abelian twisted quantum doubles for example [66]. In particular, the existing abstract constructions [12, 53] of these boundaries and defects may not be experimentally efficient in certain cases. For example, while the bulk of each phase separately has a representation in terms of familiar Pauli stabilizer models, it is unclear if the gapped boundaries and defects can also have a Pauli stabilizer description. Answering this question has practical consequences for implementation as well as for assessing the computing power of these codes. A simple systematic method to construct these boundaries and 0D defects within the Pauli Stabilizers is still lacking.
@@
In this work, we show that boundaries, domain walls and point-like defects of Abelian twisted quantum doubles can be written in terms of generalized Pauli stabilizers, thus generalizing results of Ref. 66. By providing an algorithm to construct them starting from the bulk theory, we prove this constructively. Our construction relies on local condensation, so it does not introduce new degrees of freedom [129, 128], non-Pauli stabilizers [12, 52, 53], nor rely on lattice defects [21]. It also stays within the framework of topological phases, without reverting to gauging symmetries of their underlying symmetry-protected phases [129]. It also automatically produces topologically complete models, meaning that all logical operators are macroscopic [29, 130]. Even other choices for the initial quantum double stabilizers, like spatially disconnected stabilizers [130], will automatically give models with topological order. As such, the resulting Hamiltonian for the boundary will also be described by local stabilizers. This construction offers both a simple, intuitive understanding of the boundaries as well as an efficient method that can automate boundary construction for various phases and lattices. Importantly, by relying on anyon condensation as opposed to consistency relations [93, 91, 12, 53], it extends the systematic construction of Pauli stabilizers to also include the $0D$ defects [11] of Abelian twisted quantum doubles, which is, to our knowledge, unique to this work.
@@
Our construction further facilitates the exploration of a new family of codes that comprises generalized holes, boundaries, and defects of higher-dimensional codes. For simplicity, we exemplify this new family of codes by studying the smallest composite number $d=p\times q$ where $p=q=2$. This could be realized by four-dimensional qudits or, alternatively, in the near term, by grouping qubits in pairs. We find new qualitative properties for higher-dimensional codes, such as tunable error biases, which can be exploited to simplify noise models [174]. Their syndrome extraction process can also exhibit advantages compared to the qubit case.
@@ II Summary of main results
This section summarizes the main results of and outlines our work. Readers primarily interested in explicit code constructions, decoding algorithms, and physical examples may focus on the background formalism in Sec. III and new results in Sec. IV. Readers seeking a general and systematic framework for constructing gapped boundaries, domain walls, and defects within a Pauli stabilizer formalism may consult Sec. V. The global properties of the resulting codes, including ground-state degeneracy and logical operator structure, are analyzed in Sec. VI using both microscopic and macroscopic approaches.
Sec. III describes the background and quantum double and twisted quantum double formalisms used throughout. Specifically, we work with the double of $\mathbb{Z}_{4}$, $D(\mathbb{Z}_{4})$, as a base model. We discuss periodic and open boundary conditions. In addition to the smooth and rough boundaries, we introduce an even boundary that is distinct and leads to different types of planar codes. Further, we use our boundary construction algorithm to derive a set of domain walls which generalize Bombin’s [21] $\mathbb{Z}_{2}$ twist. Sec. III.2 then moves on to describe the bulk and boundaries of the Double Semion phase model, which may be obtained by boson condensing the parent $\mathbb{Z}_{4}$ model. Throughout, we provide examples of how different combinations of boundaries and domain walls (co-dimension 1 defects) realize logical codes. For example, codes of logical dimension $D=4^{M}$ may be constructed via planar or spherical ($M=0$), open boundary condition planar codes ($M=1$), surface or toric ($M=2$), or higher $2M$-boundary surface or $M$-torus or other combinations involving twists ($M>2$). Interestingly, even boundaries lead to qubit (rather than logical $d=4$-qudits) codes. Taken together, we have codes with logical dimensions $D=4^{M}2^{M^{\prime}}$.
Building on these constructions and intuition, in Sec. IV we go beyond codes with a single bulk phase. The main tool to do so, generalizing the notion of punctures, is condensing local DS patches (co-dimension 0 defects as opposed to global condensation [66]), from the parent $\mathbb{Z}_{4}$ phase. In this composite and non-translationally invariant code, we find that $N+1$ DS patches realize an $N$ logical qubit code. Combining this result with deformations of the parent $\mathbb{Z}_{4}$ model, we get a total composite logical dimension of $D=4^{M}\times 2^{N+M^{\prime}}$. We provide the logical Pauli operators that act on the patch codes’ qubit topological codespace.
After a brief review of stabilizer codes and the noise models for qudit-based codes, we then detail how such codes can be decoded. This involves a variety of modified minimum-weight-perfect-matching decoder (MWPM), an integer linear programming decoder (ILP), and a belief propagation with ordered statistics and combination sweep decoder (BP-OSD-CS). We then use these decoders to find code capacity thresholds for the five codes we considered.
@@
After describing our new codes and their decoding strategies, Sec. V presents rigorous constructions and the main systematic methodologythath led to our codes’ discovery. Specifically, we provide an algorithm (2) that concretely illustrates how to gauge from a parent code—meaning to insert boundaries, domain walls, and 0D defects into Abelian twisted quantum doubles—to the new composite codes. We emphasize the concept of anyon condensation, which is used to conceptually understand high-level code properties, such as their interfaces. In addition, condensation results in certain code gauging strategies. These tools can be used to alter codes from the bottom up (i.e., microscopic $\rightarrow$ macroscopic) within a generalized Pauli stabilizer formalism. Importantly, we carefully examine how condensation depends on both pre-existing bulk and boundary degrees of freedom.
@@
To provide intuition, and to see the algorithm in action, a series of well-known codes and their boundaries are re-derived within this formalism in Sec. V.3. To rigorously describe our DS patch codes, we then systematically constructed boundaries and domain walls in Secs. V.4 and V.5 respectively. As an additional illustrative example, highlighting further possible modifications, the generalization from the $\mathbb{Z}_{2}$ twist to $\mathbb{Z}_{4}$ is exposed in Ex. V.4. An important example of the domain wall between $\mathbb{Z}_{4}$ and DS is treated heuristically in Ex. V.5 and, using the full machinery of our construction, in App. A. Further examples, such as the construction of 0D defects, are detailed in App. B.
@@
To insert these topological components into different topological codes under a unified framework, Sec. VI computes the ground state degeneracies in the presence of multiple boundaries, domain walls, and defects. Based on microscopic commuting projector stabilizer models, these calculations rely on counting arguments that take into account the order of old and new constraints and the changing dependencies on one another. They are central to deriving the logical dimension of our family of new quantum error correcting codes. Ultimately, the properties of the topological codes do not depend on the exact layout, which enables us to then analyze codes at scale.
@@
In Sec. VI.2, a generic code description that is agnostic to concrete stabilizer models is used to compute the codespace dimension and provide intuition about the logical operators of different codes. This is carried out using the pants decomposition of any g-genus 2D orientable manifold with the topological components of codes. This confirms the results of Sec. VI and illustrates how interested readers may further modify or generalize our codes as desired. Finally, Sec. VII concludes by discussing this work’s implications, open questions, and future research directions.
III Background
@@
The bulk of a TQFT $A$ is described by a Hamiltonian, which we denote as $H_{A}$. Generically, the TQFT is spatially supported on a surface $\mathcal{M}_{A}$ which may have a boundary $\partial\mathcal{M}_{A}$. In addition to the bulk interactions, the boundary will be described by $H_{\partial A}$ living on $\partial\mathcal{M}_{A}$. Further, two distinct TQFTs $A,B$ supported on surfaces $\partial\mathcal{M}_{A}$ and $\partial\mathcal{M}_{B}$ must have a domain wall between them if $\partial\mathcal{M}_{A}\cap\partial\mathcal{M}_{B}\neq\varnothing$. This domain wall will be described by a Hamiltonian $H_{A\cap B}$ which lives on the common boundary of the two TQFTs $\partial\mathcal{M}_{A}\cap\partial\mathcal{M}_{B}$. Altogether, the composite system will be described by a Hamiltonian.
$H_{\text{tot}}$ $=\sum_{\mathcal{M}_{A}}H_{A}+\sum_{\mathcal{M}_{B}}H_{B}+\sum_{\partial{\mathcal{M}_{A}}}H_{\partial A}+\sum_{\partial{\mathcal{M}_{B}}}H_{\partial B}$ (1)
$\quad+\sum_{\partial\mathcal{M}_{A}\cap\mathcal{M}_{B}}H_{A\cap B}$
It is evident that the boundaries of a TQFT are trivial specific examples of the domain walls between this TQFT and the vacuum TQFT. Conversely, gapped domain walls between two topologically non-trivial phases are closely related to their boundaries with the vacuum. In fact, as is discussed in Sec. V.5, a gapped domain wall between two topological phases $A$ and $B$ is precisely the gapped boundary of the topological order $A\otimes B$ [12, 114].
In the lattice picture, the models are defined on a 2D orientable surface $\mathcal{M}$, possibly containing a spatial boundary $\partial\mathcal{M}$. The surface is then triangulated into a lattice $\Gamma$ consisting of vertices $V$, edges $E$, and plaquettes $P$ (faces). Each spatial boundary $\partial\mathcal{M}$ has to be associated with a consistent boundary theory that is compatible with the bulk theory [107, 108, 52, 53, 112].
In (twisted) quantum double models [107, 92], a qudit, taking values in a finite group $G$, is associated with each edge of the lattice $e\in E$ [1]. These qudits live in a Hilbert space $\mathcal{H}_{e}$ spanned by the orthonormal basis generated by the group elements $\{|g\rangle_{e}:g\in G\}$. The total Hilbert space is then the tensor product of all the edge Hilbert spaces $\mathcal{H}_{tot}=\bigotimes_{e\in E}\mathcal{H}_{e}$. The case where the group is $\mathbb{Z}_{2}$ corresponds to the toric code, where a qubit is associated with each edge [107]. Since the basic results in this paper can also be shown in Abelian groups, we focused on the case of the cyclic group $\mathbb{Z}_{4}$, which has a simple implementation using two qubits or $d=4$ qudits (hereafter referred to as qudits). In addition, the Hamiltonian terms are built from the generalized Pauli stabilizers [75], defined below.
@@
With appropriate definitions, all physical properties of the model do not depend on the microscopic details of the lattice. Any triangulation of the surface $\mathcal{M}$ will give rise to the same topological phase [107, 92]. For simplicity, we follow the most prominent conventions, as they appear in the literature, and work with a square lattice.
In Sec. III.1 we review the $\mathbb{Z}_{4}$ surface code, which is the generalization of the surface code for the case of qudits [107]. In addition, by deriving the gapped boundaries and domain walls, we use $\mathbb{Z}_{4}$ as a tool to review defects and edges. Afterward, in Sec. III.2 we review a Doubled Semion code construction which is a twisted quantum double version of the $\mathbb{Z}_{2}$ toric code [92], and is also closely related to the $\mathbb{Z}_{4}$ code [66].
III.1 The $D(\mathbb{Z}_{4})$ Quantum Double